Need help with finding the mass of Cl for #1 and the % Chloride by mass for #2
Data Table 1. Precipitation of Chloride lon Quantity Trial 1 Trial 2 Trial 3 0i370 0.352 30,0 30.O mass ot crvble i387 0.400 0.383 C)
media%2F7a1%2F7a1f4860-1680-4aef-9c8c-7a
media%2Fb70%2Fb7036a19-c377-4785-aa90-b2
Answer #1

1.

Trial (1)

Mass of AgCl = 0.387 g.

Molar mass of AgCl = 107.87 g/mol

Amount of AgCl = mass / molar mass = 0.387 / 107.87 = 0.00359 mol

From the fromula,

1 mol AgCl has 1 mol Cl

then, 0.00359 mol AgCl has 0.00359 mol Cl

Therefore,

Mass of Cl = moles * molar mass = 0.00359 * 35.45 = 0.127 g.

2.

Trial (1)

% by mass of Cl = Mass of Cl * 100 / Mass of unknown sample = 0.127 * 100 / 0.354 = 35.9 %